An apple is falling down from a height of 20m and then hit a


person. Assume that the mass of the apple is 200g. The


duration of impact is 0.001 sec and that the area of the


bone is 1 cm(square). (a) Calculate the impulsive force applied to


the person. (b) If the rupture strength for body tissue is


about 5 X 1000000dyn/cm(sqare), will the impact force lead to serious


injury for the person?

1
Expert's answer
2021-11-14T17:13:53-0500

a)

"F=\frac pt=\frac {mv}t=\frac mt\sqrt{2gh}=4000~N,"

b)

"\frac FS=4000>\sigma =50,"

the impact force will lead to serious injury for the person.